OpenStudy (anonymous):

3 √52 +2√208 – 2 √117

OpenStudy (anonymous):

you would simplify radical 52 to radical 4 and radical 13 right?

OpenStudy (mathstudent55):

Factor 52, 208, and 117 to see which perfect squares can be facored out. 52 = 4 x 13, so\[3 \sqrt{52} = 3 \sqrt{4 \times 13} = 3 \sqrt{4} \sqrt{13} = 6 \sqrt{13}\] Try the same with 208 and 117

OpenStudy (mathstudent55):

Yes, you are correct, radical 52 is radical 4 times radical 13. Dothe same simplification to radical 208 and radical 117

OpenStudy (anonymous):

I need help with that part

OpenStudy (mathstudent55):

Since you know that radical 52 includes 13 in it, try dividing both 208 and 117 by 13. You see that radical 208 = radical 16 times radical 13, and radical 117 is radical 9 times radical 13

OpenStudy (anonymous):

so 6 radical 13 +4 radical 13 -3 radical 13

OpenStudy (mathstudent55):

\[3 \sqrt{52} + 2 \sqrt{208} - 2 \sqrt{117}\] \[3 \sqrt{4} \sqrt{13} + 2 \sqrt{16} \sqrt{13} -2 \sqrt{9} \sqrt{13}\]

OpenStudy (anonymous):

\[6\sqrt{13} +4\sqrt{13} -3\sqrt{13}\]

OpenStudy (mathstudent55):

Remember that radical 208 and radical 117 have numbers multiplying them

OpenStudy (anonymous):

sorry saw my mistake

OpenStudy (mathstudent55):

Great, you're getting there

OpenStudy (anonymous):

Ya im done with this prob. YAy